OpenStudy (anonymous):

given f(x) shown below, complete the equation for the inverse of f(x). if necessary use the slash mark (/) for the division problem f(x)=3x+6 F^-1(y)=_____

OpenStudy (anonymous):

i really need help

OpenStudy (luigi0210):

To find the inverse all you need to do is switch y and x around.

OpenStudy (anonymous):

yes i know that but i forgot because people were telling me different things and i got all of my ?s wronggg

OpenStudy (luigi0210):

Well inverse is just switching. So f(x)=3x+6 becomes x=3y+6 Now just solve for y

OpenStudy (anonymous):

-6 to both sides right

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

y=1/3?

OpenStudy (debbieg):

x=3y+6 Subtracting 6 from both sides should give you x-6=3y Then solve for y. You will get y as an expression of x, a FUNCTION, not a NUMBER.

OpenStudy (anonymous):

\[f ^{-1}(y)=\frac{ 1 }{ 3 }\]

OpenStudy (debbieg):

But @romanortiz65 , as some of us were telling you yesterday, your teacher's notation is very weird because in your question you have: F^-1(y)= Usually when you find an inverse, you give it in terms of x, not in terms of y. So the way that @Luigi0210 described above is absolutely the standard approach to this type of problem (switch x and y in the original function and then solve that expression for y), but your teacher asking for the function to be in terms of y, which you would get if you DON'T switch the variables and simply solve the original equation for x.

OpenStudy (anonymous):

\[f ^{-1}(y)=\]

OpenStudy (debbieg):

You have \[\large x=3y+6\] you subtract 6 from each side and get \[\large x-6=3y\] How are you getting from that, to \(y=\dfrac{1}{3}\)??

OpenStudy (anonymous):

it should be written as this

OpenStudy (anonymous):

ohhhhh i see

OpenStudy (anonymous):

\[\frac{ x-6 }{ 3 }\]

OpenStudy (anonymous):

is the answer

OpenStudy (debbieg):

Yes, after you solve the equation for y as a function of x, you TYPICALLY write \[y=f ^{-1}(x)\] This is the oddity of the notation that you are being asked for, since you are giving a function of y it would be \[x=f ^{-1}(y)\] and I'm not sure what to make of that, lol.

OpenStudy (debbieg):

Yes, (x-6)/3 is correct.

OpenStudy (anonymous):

ok no wonder i got all my test wrong i kept on writting them as x-6/3